OpenStudy (anonymous):

what is the inverse of f(x)=3-2x

zepdrix (zepdrix):

\(\Large\rm f(x)=3-2x\) Hey Alisha :)

zepdrix (zepdrix):

Step 1: Replace f(x) with y, \(\Large\rm y=3-2x\)

zepdrix (zepdrix):

Step 2: Swap your x's and y's, \(\Large\rm x=3-2y\)

zepdrix (zepdrix):

And then solve for this y!

zepdrix (zepdrix):

Hmmm so what should we do? Maybe subtract 3 from each side? \(\Large\rm x-3=-2y\) What else, what do you think? :)